# Hui Xiong2, Bin Yang1

### v1

@@ Appendix C Full Experimental Results
Dataset Metric CATCH Modern iTrans DualTF ATrans DC TsNet Patch DLin NLin AE Ocsvm IF PCA HBOS
CICIDS R-R 0.786 0.277 0.494 0.586 0.243 0.283 0.369 0.389 0.523 0.386 0.000 0.986 0.057 0.394 0.009
R-P 0.004 0.002 0.002 0.001 0.001 0.001 0.002 0.002 0.003 0.001 0.000 0.004 0.001 0.001 0.000
R-F 0.007 0.003 0.004 0.003 0.003 0.002 0.003 0.004 0.005 0.003 0.000 0.008 0.001 0.003 0.000
CalIt2 R-R 0.278 0.318 0.344 0.209 0.161 0.205 0.161 0.424 0.315 0.325 0.055 0.359 0.031 0.280 0.309
R-P 0.139 0.069 0.113 0.056 0.105 0.031 0.095 0.077 0.091 0.064 0.067 0.096 0.125 0.099 0.086
R-F 0.185 0.114 0.170 0.088 0.127 0.053 0.120 0.131 0.142 0.107 0.060 0.152 0.049 0.147 0.134
Credit R-R 0.735 0.716 0.828 0.564 0.150 0.234 0.716 0.725 0.701 0.721 0.135 0.892 0.292 0.709 0.907
R-P 0.054 0.052 0.024 0.002 0.003 0.002 0.051 0.052 0.049 0.052 0.047 0.012 0.156 0.008 0.009
R-F 0.101 0.097 0.047 0.004 0.007 0.003 0.096 0.096 0.092 0.097 0.070 0.023 0.204 0.015 0.017
GECCO R-R 0.795 0.644 0.266 0.146 0.063 0.040 0.782 0.366 0.790 0.274 0.188 0.953 0.030 0.986 0.361
R-P 0.065 0.086 0.111 0.039 0.013 0.008 0.053 0.289 0.042 0.296 0.021 0.006 0.214 0.041 0.016
R-F 0.119 0.152 0.156 0.062 0.022 0.013 0.099 0.323 0.080 0.285 0.038 0.011 0.052 0.079 0.031
Genesis R-R 0.497 0.174 0.507 0.550 0.855 0.079 0.211 0.385 0.180 0.162 0.356 1.000 0.077 0.325 0.861
R-P 0.119 0.014 0.057 0.030 0.005 0.016 0.086 0.067 0.059 0.013 0.047 0.003 0.006 0.035 0.002
R-F 0.192 0.026 0.102 0.058 0.010 0.027 0.122 0.114 0.089 0.024 0.083 0.005 0.011 0.063 0.005
MSL R-R 0.241 0.194 0.182 0.049 0.150 0.126 0.224 0.176 0.202 0.162 0.110 0.202 0.112 0.237 0.199
R-P 0.150 0.129 0.125 0.267 0.143 0.179 0.130 0.136 0.130 0.137 0.219 0.098 0.180 0.113 0.119
R-F 0.185 0.155 0.148 0.084 0.146 0.148 0.164 0.153 0.158 0.149 0.146 0.132 0.138 0.153 0.149
NYC R-R 0.208 0.000 0.223 0.208 0.208 0.208 0.000 0.000 0.000 0.000 0.208 0.461 0.208 0.289 0.000
R-P 1.000 0.000 0.023 0.143 0.333 0.250 0.000 0.000 0.000 0.000 0.059 0.040 0.056 0.053 0.000
R-F 0.344 0.000 0.042 0.169 0.256 0.227 0.000 0.000 0.000 0.000 0.092 0.074 0.088 0.089 0.000
PSM R-R 0.450 0.376 0.470 0.410 0.133 0.211 0.455 0.370 0.399 0.359 0.230 0.133 0.139 0.505 0.265
R-P 0.557 0.553 0.533 0.456 0.374 0.274 0.537 0.557 0.584 0.587 0.444 0.385 0.802 0.459 0.467
R-F 0.498 0.448 0.499 0.432 0.197 0.238 0.493 0.444 0.474 0.446 0.303 0.198 0.237 0.481 0.338
SMD R-R 0.478 0.378 0.360 0.187 0.126 0.276 0.385 0.426 0.400 0.384 0.079 0.323 0.122 0.404 0.297
R-P 0.095 0.092 0.115 0.055 0.101 0.042 0.110 0.121 0.124 0.131 0.706 0.067 0.454 0.123 0.062
R-F 0.158 0.148 0.175 0.085 0.112 0.073 0.171 0.189 0.189 0.196 0.143 0.111 0.192 0.188 0.103
Contextual4.9 R-R 0.733 0.341 0.697 0.232 0.347 0.361 0.556 0.709 0.656 0.659 0.603 0.559 0.396 0.181 0.053
R-P 0.586 0.071 0.559 0.041 0.057 0.047 0.139 0.267 0.440 0.125 0.670 0.240 0.577 0.061 0.025
R-F 0.651 0.118 0.620 0.069 0.097 0.083 0.222 0.388 0.526 0.210 0.635 0.336 0.470 0.092 0.034
Contextual7.2 R-R 0.692 0.335 0.725 0.453 0.291 0.345 0.597 0.697 0.717 0.654 0.580 0.553 0.360 0.192 0.061
R-P 0.744 0.100 0.546 0.070 0.080 0.060 0.227 0.353 0.365 0.164 0.775 0.304 0.633 0.087 0.053
R-F 0.717 0.154 0.623 0.122 0.125 0.103 0.328 0.468 0.483 0.263 0.664 0.392 0.459 0.120 0.057
Global4.8 R-R 0.919 0.726 0.874 0.206 0.859 0.327 0.855 0.887 0.870 0.823 0.887 0.996 0.759 0.733 0.178
R-P 0.939 0.292 1.000 0.073 0.041 0.043 0.900 1.000 0.978 0.362 0.747 0.338 0.766 0.113 0.116
R-F 0.929 0.416 0.933 0.108 0.079 0.076 0.877 0.940 0.921 0.503 0.811 0.505 0.762 0.196 0.140
Global7.2 R-R 0.922 0.716 0.881 0.742 0.840 0.320 0.832 0.916 0.869 0.801 0.849 0.982 0.761 0.712 0.083
R-P 0.832 0.297 0.773 0.078 0.052 0.072 0.984 0.874 0.917 0.460 0.871 0.457 0.834 0.125 0.070
R-F 0.874 0.420 0.824 0.141 0.097 0.117 0.902 0.894 0.892 0.584 0.860 0.624 0.796 0.213 0.076
Seasonal4.8 R-R 0.475 0.244 0.464 0.247 0.238 0.170 0.453 0.441 0.433 0.364 0.280 0.478 0.280 0.238 0.187
R-P 0.999 0.062 1.000 0.597 0.864 0.500 1.000 1.000 1.000 0.966 0.630 0.113 0.610 0.065 0.041
R-F 0.644 0.099 0.634 0.350 0.373 0.254 0.624 0.612 0.604 0.529 0.387 0.182 0.383 0.102 0.067
Seasonal7.7 R-R 0.480 0.254 0.425 0.089 0.173 0.191 0.423 0.409 0.397 0.355 0.278 0.453 0.278 0.187 0.223
R-P 0.944 0.071 1.000 0.112 0.962 0.588 0.993 0.917 1.000 0.926 0.692 0.178 0.766 0.063 0.149
R-F 0.636 0.111 0.596 0.099 0.293 0.289 0.594 0.566 0.569 0.513 0.396 0.256 0.408 0.094 0.179
Shapelet4.9 R-R 0.447 0.297 0.431 0.057 0.154 0.174 0.398 0.445 0.445 0.249 0.279 0.496 0.244 0.224 0.116
R-P 0.944 0.065 0.679 0.104 0.433 0.382 0.964 0.348 0.778 0.213 0.479 0.083 0.542 0.057 0.036
R-F 0.607 0.107 0.527 0.074 0.227 0.240 0.564 0.391 0.566 0.230 0.353 0.143 0.336 0.091 0.055
Shapelet7.4 R-R 0.471 0.304 0.473 0.197 0.138 0.343 0.442 0.457 0.466 0.365 0.277 0.517 0.253 0.315 0.096
R-P 0.980 0.083 0.620 0.123 0.309 0.153 0.818 0.479 0.598 0.211 0.583 0.118 0.489 0.114 0.042
R-F 0.636 0.130 0.536 0.151 0.191 0.212 0.574 0.468 0.524 0.268 0.375 0.193 0.333 0.168 0.058
Mixture5.7 R-R 0.601 0.341 0.508 0.347 0.076 0.119 0.515 0.487 0.705 0.364 0.279 0.954 0.279 0.954 0.718
R-P 0.455 0.091 0.355 0.062 0.100 0.333 0.509 0.344 0.066 0.233 0.066 0.053 0.627 0.091 0.096
R-F 0.518 0.143 0.418 0.106 0.086 0.175 0.512 0.403 0.121 0.284 0.107 0.100 0.386 0.165 0.170
Mixture5.9 R-R 0.670 0.293 0.595 0.343 0.057 0.145 0.599 0.558 0.594 0.465 0.279 0.998 0.267 1.000 0.806
R-P 0.516 0.148 0.401 0.089 0.070 0.443 0.511 0.342 0.108 0.217 0.092 0.000 0.704 0.000 0.116
R-F 0.583 0.197 0.479 0.141 0.063 0.218 0.551 0.424 0.183 0.296 0.139 0.000 0.387 0.000 0.203
Trend4.8 R-R 0.910 0.492 0.562 0.277 0.023 0.082 0.474 0.483 0.625 0.538 0.280 0.997 0.268 1.000 0.799
R-P 0.247 0.157 0.205 0.073 0.037 0.167 0.273 0.169 0.064 0.117 0.048 0.000 0.600 0.002 0.050
R-F 0.388 0.238 0.300 0.115 0.029 0.110 0.347 0.251 0.116 0.192 0.081 0.001 0.371 0.004 0.094
Trend7.8 R-R 0.529 0.680 0.561 0.316 0.095 0.015 0.681 0.546 0.715 0.432 0.279 0.996 0.220 1.000 0.768
R-P 0.242 0.122 0.168 0.099 0.146 0.048 0.187 0.165 0.082 0.155 0.076 0.001 0.641 0.002 0.064
R-F 0.332 0.207 0.258 0.150 0.115 0.023 0.294 0.253 0.147 0.229 0.119 0.002 0.327 0.004 0.117
$1^{st}$ Count 22 0 5 1 0 0 3 3 0 2 3 11 7 5 1

### v3

@@ Appendix C Full Experimental Results
Dataset Metric CATCH Modern iTrans DualTF ATrans DC TsNet Patch DLin NLin AE TFAD Ocsvm IF PCA HBOS
CICIDS R-R 0.786 0.277 0.494 0.586 0.243 0.283 0.369 0.389 0.523 0.386 0.000 0.043 0.986 0.057 0.394 0.009
R-P 0.004 0.002 0.002 0.001 0.001 0.001 0.002 0.002 0.003 0.001 0.000 0.001 0.004 0.001 0.001 0.000
R-F 0.007 0.003 0.004 0.003 0.003 0.002 0.003 0.004 0.005 0.003 0.000 0.001 0.008 0.001 0.003 0.000
CalIt2 R-R 0.278 0.318 0.344 0.209 0.161 0.205 0.161 0.424 0.315 0.325 0.055 0.074 0.359 0.031 0.280 0.309
R-P 0.139 0.069 0.113 0.056 0.105 0.031 0.095 0.077 0.091 0.064 0.067 0.038 0.096 0.125 0.099 0.086
R-F 0.185 0.114 0.170 0.088 0.127 0.053 0.120 0.131 0.142 0.107 0.060 0.051 0.152 0.049 0.147 0.134
Credit R-R 0.735 0.716 0.828 0.564 0.150 0.234 0.716 0.725 0.701 0.721 0.135 0.020 0.892 0.292 0.709 0.907
R-P 0.054 0.052 0.024 0.002 0.003 0.002 0.051 0.052 0.049 0.052 0.047 0.005 0.012 0.156 0.008 0.009
R-F 0.101 0.097 0.047 0.004 0.007 0.003 0.096 0.096 0.092 0.097 0.070 0.007 0.023 0.204 0.015 0.017
GECCO R-R 0.795 0.644 0.266 0.146 0.063 0.040 0.782 0.366 0.790 0.274 0.188 0.030 0.953 0.030 0.986 0.361
R-P 0.065 0.086 0.111 0.039 0.013 0.008 0.053 0.289 0.042 0.296 0.021 0.009 0.006 0.214 0.041 0.016
R-F 0.119 0.152 0.156 0.062 0.022 0.013 0.099 0.323 0.080 0.285 0.038 0.013 0.011 0.052 0.079 0.031
Genesis R-R 0.497 0.174 0.507 0.550 0.855 0.079 0.211 0.385 0.180 0.162 0.356 0.000 1.000 0.077 0.325 0.861
R-P 0.119 0.014 0.057 0.030 0.005 0.016 0.086 0.067 0.059 0.013 0.047 0.000 0.003 0.006 0.035 0.002
R-F 0.192 0.026 0.102 0.058 0.010 0.027 0.122 0.114 0.089 0.024 0.083 0.000 0.005 0.011 0.063 0.005
MSL R-R 0.241 0.194 0.182 0.049 0.150 0.126 0.224 0.176 0.202 0.162 0.110 0.138 0.202 0.112 0.237 0.199
R-P 0.150 0.129 0.125 0.267 0.143 0.179 0.130 0.136 0.130 0.137 0.219 0.107 0.098 0.180 0.113 0.119
R-F 0.185 0.155 0.148 0.084 0.146 0.148 0.164 0.153 0.158 0.149 0.146 0.120 0.132 0.138 0.153 0.149
NYC R-R 0.208 0.000 0.223 0.208 0.208 0.208 0.000 0.000 0.000 0.000 0.208 0.000 0.461 0.208 0.289 0.000
R-P 1.000 0.000 0.023 0.143 0.333 0.250 0.000 0.000 0.000 0.000 0.059 0.000 0.040 0.056 0.053 0.000
R-F 0.344 0.000 0.042 0.169 0.256 0.227 0.000 0.000 0.000 0.000 0.092 0.000 0.074 0.088 0.089 0.000
PSM R-R 0.450 0.376 0.470 0.410 0.133 0.211 0.455 0.370 0.399 0.359 0.230 0.054 0.133 0.139 0.505 0.265
R-P 0.557 0.553 0.533 0.456 0.374 0.274 0.537 0.557 0.584 0.587 0.444 0.282 0.385 0.802 0.459 0.467
R-F 0.498 0.448 0.499 0.432 0.197 0.238 0.493 0.444 0.474 0.446 0.303 0.091 0.198 0.237 0.481 0.338
SMD R-R 0.478 0.378 0.360 0.187 0.126 0.276 0.385 0.426 0.400 0.384 0.079 0.053 0.323 0.122 0.404 0.297
R-P 0.095 0.092 0.115 0.055 0.101 0.042 0.110 0.121 0.124 0.131 0.706 0.043 0.067 0.454 0.123 0.062
R-F 0.158 0.148 0.175 0.085 0.112 0.073 0.171 0.189 0.189 0.196 0.143 0.047 0.111 0.192 0.188 0.103
ASD R-R 0.303 0.255 0.291 0.145 0.097 0.116 0.328 0.265 0.232 0.215 0.174 0.062 0.166 0.196 0.312 0.217
R-P 0.212 0.185 0.165 0.165 0.081 0.070 0.218 0.182 0.232 0.181 0.222 0.077 0.087 0.455 0.106 0.084
R-F 0.219 0.192 0.176 0.098 0.084 0.083 0.237 0.185 0.207 0.179 0.185 0.066 0.102 0.266 0.146 0.111
Contextual4.9 R-R 0.733 0.341 0.697 0.232 0.347 0.361 0.556 0.709 0.656 0.659 0.603 0.070 0.559 0.396 0.181 0.053
R-P 0.586 0.071 0.559 0.041 0.057 0.047 0.139 0.267 0.440 0.125 0.670 0.041 0.240 0.577 0.061 0.025
R-F 0.651 0.118 0.620 0.069 0.097 0.083 0.222 0.388 0.526 0.210 0.635 0.052 0.336 0.470 0.092 0.034
Contextual7.2 R-R 0.692 0.335 0.725 0.453 0.291 0.345 0.597 0.697 0.717 0.654 0.580 0.096 0.553 0.360 0.192 0.061
R-P 0.744 0.100 0.546 0.070 0.080 0.060 0.227 0.353 0.365 0.164 0.775 0.084 0.304 0.633 0.087 0.053
R-F 0.717 0.154 0.623 0.122 0.125 0.103 0.328 0.468 0.483 0.263 0.664 0.090 0.392 0.459 0.120 0.057
Global4.8 R-R 0.919 0.726 0.874 0.206 0.859 0.327 0.855 0.887 0.870 0.823 0.887 0.078 0.996 0.759 0.733 0.178
R-P 0.939 0.292 1.000 0.073 0.041 0.043 0.900 1.000 0.978 0.362 0.747 0.046 0.338 0.766 0.113 0.116
R-F 0.929 0.416 0.933 0.108 0.079 0.076 0.877 0.940 0.921 0.503 0.811 0.058 0.505 0.762 0.196 0.140
Global7.2 R-R 0.922 0.716 0.881 0.742 0.840 0.320 0.832 0.916 0.869 0.801 0.849 0.068 0.982 0.761 0.712 0.083
R-P 0.832 0.297 0.773 0.078 0.052 0.072 0.984 0.874 0.917 0.460 0.871 0.060 0.457 0.834 0.125 0.070
R-F 0.874 0.420 0.824 0.141 0.097 0.117 0.902 0.894 0.892 0.584 0.860 0.064 0.624 0.796 0.213 0.076
Seasonal4.8 R-R 0.475 0.244 0.464 0.247 0.238 0.170 0.453 0.441 0.433 0.364 0.280 0.000 0.478 0.280 0.238 0.187
R-P 0.999 0.062 1.000 0.597 0.864 0.500 1.000 1.000 1.000 0.966 0.630 0.000 0.113 0.610 0.065 0.041
R-F 0.644 0.099 0.634 0.350 0.373 0.254 0.624 0.612 0.604 0.529 0.387 0.000 0.182 0.383 0.102 0.067
Seasonal7.7 R-R 0.480 0.254 0.425 0.089 0.173 0.191 0.423 0.409 0.397 0.355 0.278 0.232 0.453 0.278 0.187 0.223
R-P 0.944 0.071 1.000 0.112 0.962 0.588 0.993 0.917 1.000 0.926 0.692 0.323 0.178 0.766 0.063 0.149
R-F 0.636 0.111 0.596 0.099 0.293 0.289 0.594 0.566 0.569 0.513 0.396 0.270 0.256 0.408 0.094 0.179
Shapelet4.9 R-R 0.447 0.297 0.431 0.057 0.154 0.174 0.398 0.445 0.445 0.249 0.279 0.012 0.496 0.244 0.224 0.116
R-P 0.944 0.065 0.679 0.104 0.433 0.382 0.964 0.348 0.778 0.213 0.479 0.043 0.083 0.542 0.057 0.036
R-F 0.607 0.107 0.527 0.074 0.227 0.240 0.564 0.391 0.566 0.230 0.353 0.018 0.143 0.336 0.091 0.055
Shapelet7.4 R-R 0.471 0.304 0.473 0.197 0.138 0.343 0.442 0.457 0.466 0.365 0.277 0.221 0.517 0.253 0.315 0.096
R-P 0.980 0.083 0.620 0.123 0.309 0.153 0.818 0.479 0.598 0.211 0.583 0.235 0.118 0.489 0.114 0.042
R-F 0.636 0.130 0.536 0.151 0.191 0.212 0.574 0.468 0.524 0.268 0.375 0.228 0.193 0.333 0.168 0.058
Mixture5.7 R-R 0.601 0.341 0.508 0.347 0.076 0.119 0.515 0.487 0.705 0.364 0.279 0.159 0.954 0.279 0.954 0.718
R-P 0.455 0.091 0.355 0.062 0.100 0.333 0.509 0.344 0.066 0.233 0.066 0.113 0.053 0.627 0.091 0.096
R-F 0.518 0.143 0.418 0.106 0.086 0.175 0.512 0.403 0.121 0.284 0.107 0.132 0.100 0.386 0.165 0.170
Mixture8.9 R-R 0.670 0.293 0.595 0.343 0.057 0.145 0.599 0.558 0.594 0.465 0.279 0.178 0.998 0.267 1.000 0.806
R-P 0.516 0.148 0.401 0.089 0.070 0.443 0.511 0.342 0.108 0.217 0.092 0.089 0.000 0.704 0.000 0.116
R-F 0.583 0.197 0.479 0.141 0.063 0.218 0.551 0.424 0.183 0.296 0.139 0.118 0.000 0.387 0.000 0.203
Trend4.8 R-R 0.910 0.492 0.562 0.277 0.023 0.082 0.474 0.483 0.625 0.538 0.280 0.000 0.997 0.268 1.000 0.799
R-P 0.247 0.157 0.205 0.073 0.037 0.167 0.273 0.169 0.064 0.117 0.048 0.000 0.000 0.600 0.002 0.050
R-F 0.388 0.238 0.300 0.115 0.029 0.110 0.347 0.251 0.116 0.192 0.081 0.000 0.001 0.371 0.004 0.094
Trend7.8 R-R 0.529 0.680 0.561 0.316 0.095 0.015 0.681 0.546 0.715 0.432 0.279 0.183 0.996 0.220 1.000 0.768
R-P 0.242 0.122 0.168 0.099 0.146 0.048 0.187 0.165 0.082 0.155 0.076 0.127 0.001 0.641 0.002 0.064
R-F 0.332 0.207 0.258 0.150 0.115 0.023 0.294 0.253 0.147 0.229 0.119 0.150 0.002 0.327 0.004 0.117
$1^{st}$ Count 23 0 5 1 0 0 4 3 0 2 3 0 11 9 7 1